A mixture of 4% disinfectant solution is to be made from 8% and 2% disinfectant solutions. How much of each solution should be used if 33 gallons of 4% solution are needed?

Answer by mananth(16946)

Percent ---------------- quantity
disinfectant I 8.00% ---------------- x gallons
disinfectant II 2.00% ------ 33 - x gallons
Mixture 4.00% ---------------- 33
Total 33 gallons
8.00% x + 2.00% ( 33 - x ) = 4.00% * 33
8 x + 2 ( 33 - x ) = 132
8 x + 66 - 2 x = 132
8 x - 2 x = 132 - -66
6 x = 66
/ 6
x = 11 gallons 8.00% disinfectant I
22 gallons 2.00% disinfectant II

m.ananth@hotmail.ca